# RL Zero: Zero-Shot Language to Behaviors without any supervision

### v1

@@ Abstract
Rewards remain an uninterpretable way to specify tasks for Reinforcement Learning, as humans are often unable to predict the optimal behavior of any given reward function, leading to poor reward design and reward hacking. Language presents an appealing way to communicate intent to agents and bypass reward design, but prior efforts to do so have been limited by costly and unscalable labeling efforts. In this work, we propose a method for a completely unsupervised alternative to grounding language instructions in a zero-shot manner to obtain policies. We present a solution that takes the form of imagine, project, and imitate: The agent imagines the observation sequence corresponding to the language description of a task, projects the imagined sequence to our target domain, and grounds it to a policy. Video-language models allow us to imagine task descriptions that leverage knowledge of tasks learned from internet-scale video-text mappings. The challenge remains to ground these generations to a policy. In this work, we show that we can achieve a zero-shot language-to-behavior policy by first grounding the imagined sequences in real observations of an unsupervised RL agent and using a closed-form solution to imitation learning that allows the RL agent to mimic the grounded observations. Our method, RLZero, is the first to our knowledge to show zero-shot language to behavior generation abilities without any supervision on a variety of tasks on simulated domains. We further show that RLZero can also generate policies zero-shot from cross-embodied videos such as those scraped from YouTube.
@@ 1 Introduction
Underlying the many successes of RL lies the engineering challenge of task specification, where a skilled expert painstakingly designs a reward function. Not only does this restrict the scaling of RL agents, but it also makes those agents uninterpretable to any user inexperienced with reward design. Even for experts, reasoning about simple reward functions is generally infeasible because these functions can be easily hacked (Krakovna, 2018; Amodei et al., 2016; Dulac-Arnold et al., 2021) to produce behaviors that do not align with human intent. Language is an expressive communication channel for human intent and allows bypassing reward design, but learning a mapping from language to behaviors has historically required collecting and annotating behaviors that correspond to language (Jang et al., 2022; O’Neill et al., 2023). This strategy is impractical at scale where samples from the agent’s large space of behaviors need to be labeled. Instead, an approach that strictly makes use of models learned in a purely unsupervised way becomes desirable.
How can generalist agents interpret language commands into behaviors? Large-scale multimodal foundation models (Wang et al., 2022) provide us with part of the solution. Trained on large amounts of internet data, they can assist in generating video segments that communicate what performing a task entails. An issue with the video generation models is that they may generate video frames demonstrating tasks that are out of distribution for the current agent’s domain; for instance, the current agent can be in a simulated environment, and the video generation models produce generations resembling the real world with potentially differing dynamics. In this work, we propose to fix this problem by projecting frames to agent observations under a semantic similarity scoring metric (Radford et al., 2021b; Zhai et al., 2023). This frame-by-frame similarity search allows us to produce a sequence of observations grounded in the agent’s interaction history and presents an expectation of what the task would look like grounded in agent’s observation space. However, the discovered frame sequence might not adhere to environment dynamics or even be feasible. This presents us with our next question: How do we generate behaviors that resemble the grounded imagined trajectories in a zero-shot [^1] Our usage of the term zero-shot is based on the zero shot RL definition from Touati et al. (2023) manner? For this, we rely on unsupervised RL techniques that allow us zero-shot inference for a diverse set of behaviors specified by any reward function.
Unsupervised RL offers an ideal tool for zero-shot behavior inference, enabling an agent to leverage task-agnostic prior interactions with the environment to encode diverse behaviors. A generalist agent should be capable of a wide variety of skills, and these skills are expected to be learned through the course of the agent’s prior interactions with the environment. By comparison, providing expert demonstrations and annotating each skill of an agent with a language description is prohibitively expensive. Classical reinforcement learning algorithms sidestep the first issue of requiring expert demonstration but still require specifying a reward function. Prior research (Rocamonde et al., 2023; Baumli et al., 2023) has used large Vision-Language Models (VLMs) to obtain rewards for a language command. Even when a reward function is provided, training a policy for a reward function from scratch each time is time-consuming, undesirable, and potentially unsafe. Instead, we want agents that are sufficiently capable of solving a wide variety of tasks in a zero-shot manner (without further training or gradient updates). Hence, we turn to Unsupervised RL techniques, sometimes termed as Behavior Foundation Models (BFM) (Touati & Ollivier, 2021; Park et al., 2024a; Agarwal et al., 2024), that allow learning behaviors for all possible reward functions subject to model capacity constraints. BFMs work by pretraining optimal policies for all rewards defined in the span of learned state features. During inference, the optimal policy corresponding to a particular reward function can be obtained in closed form.
@@
Figure 1: RLZero framework of imagine, project and imitate: A video trajectory is imagined using the text prompt in the agent’s observation space and projected to real agent observations. Using observation-only imitation learning, the generated trajectory is grounded in a policy that mimics the behavior demonstrated by the video.
@@
Our method sidesteps the requirement of reward functions and instead frames the problem of language-to-skill inference as a state-only distribution matching to the grounded imagined trajectories. Notably, this work leverages the capability of unsupervised RL methods to provide a zero-shot solution to distribution matching. Finally, our method parallels the imagination capabilities of humans to picture in their mind possibilities in the real world (Sarbin, 2004; Sarbin & Juhasz, 1970; Pylyshyn, 2002). RLZero works in three simple steps: a) Imagine: Imagine trajectories given a language command: We use generative models to visualize the desired tasks specified by text as sequences of frames. These trajectories are action-free and may not be dynamically feasible. b) Project: The frames of imagined trajectories are projected to real observations of the agent. c) Imitate: RLZero leverages the agent’s prior environmental interactions to directly output a policy that matches the state visitation distribution of the imagined trajectories. Our experiments show that RLZero is a promising approach to designing an interpretable link connecting humans to RL agents. We demonstrate that RLZero is an effective method on a variety of tasks where reward function design would require a full-time reward engineer. We show that RLZero also opens possibilities for zero-shot cross-embodiment transfer, a first approach to be able to do this to our knowledge. Our contribution is the framework of imagine, project, and zero-shot imitate, which diverges from the classical wisdom of using VLMs as reward functions—which can be hacked—and instead focusing on zero-shot imitation with unsupervised RL, which admits a unique solution that matches the imagined behavior.
@@ 2 Related Work
Language and Control: There is a rich history of using language to solve various tasks in RL: task specification (Thomason et al., 2015; Goyal et al., 2021b; Ma et al., 2023; Baumli et al., 2023; Rocamonde et al., 2023; Stepputtis et al., 2020; Brohan et al., 2022; 2023), transfer and generalization (Goyal et al., 2021a; Jang et al., 2022; Liang et al., 2023), using language to provide hierarchies that allow for solving long-horizon tasks (Ahn et al., 2022; Jiang et al., 2019), driving exploration (Goyal et al., 2019; Harrison et al., 2017; Wang et al., 2023; Ma et al., 2024), human-in-the-loop learning (Chen et al., 2020; Chevalier-Boisvert et al., 2019), giving feedback to AI agents (Wang et al., 2024b), reward design (Yu et al., 2023), etc. Most existing methods either require labels for mapping language to low-level actions or generate reward functions that need to be trained by interacting with the environment to generate a low-level control policy. Recent work (Mazzaglia et al., 2024) proposed an unsupervised approach to grounding language to low-level skills but requires re-training the RL agent for each given task prompt. In contrast, our work presents a method that allows for zero-shot mapping of languages to low-level skills. A large portion of prior work has been limited to using language in a setting where expert demonstrations are provided, but this puts a heavy burden on data collection to cover the large number of skills possible in the environment, which quickly becomes impractical considering the vast array of interactions intelligent agents can perform with their environments. Our approach forgoes this limitation by relying on a zero-shot RL agent capable of mimicking arbitrary imaginations generated for a given text.
Zero-shot RL: Zero-shot RL promises the ability to quickly produce optimal policies for any given task defined by a reward function. A wide variety of methods have been developed to achieve zero-shot RL, which are in some ways generalizations of multi-task RL (Caruana, 1997). Most of these works assume a class of tasks where they can produce policies zero-shot. These tasks can be goal-conditioned (Kaelbling, 1993; Durugkar et al., 2021; Agarwal et al., 2023; Sikchi et al., 2023; Ma et al., 2022b), a linear span of certain state-features (Dayan, 1993; Barreto et al., 2017; Blier et al., 2021b; Touati & Ollivier, 2021; Park et al., 2024a; Agarwal et al., 2024) or some combination of some skills (Eysenbach et al., 2018; 2022; Park et al., 2024b). Other methods (Agarwal et al., 2024) have looked at representing all policies in RL but learning this becomes cumbersome in high dimensions. Recent works (Wu et al., 2018; Touati & Ollivier, 2021; Touati et al., 2023; Park et al., 2024a; Agarwal et al., 2024) employ a successor measure-based representation learning objective to be able to provide near-optimal policies for arbitrary reward function subject to model capacity constraints. Our work leverages these methods and finds the best reward supported by the representations that will produce the language-conditioned imagined trajectory.
@@ 3 Preliminaries
We consider a learning agent in a Markov Decision Process (MDP) (Puterman, 2014; Sutton & Barto, 2018) which is defined as a tuple: $\mathcal{M}=({\mathcal{S}},{\mathcal{A}},p,R,\gamma,d_{0})$ where ${\mathcal{S}}$ and ${\mathcal{A}}$ denote the state and action spaces respectively, $p$ denotes the transition function with $p(s^{\prime}|s,a)$ indicating the probability of transitioning from $s$ to $s^{\prime}$ taking action $a$; $R$ denotes the reward function, $\gamma\in(0,1)$ specifies the discount factor and $d_{0}$ denotes the initial state distribution. The reinforcement learning objective is to obtain a policy $\pi:{\mathcal{S}}\rightarrow\Delta({\mathcal{A}})$ that maximizes expected return: $\mathbb{E}_{\pi}[{\sum_{t=0}^{\infty}\gamma^{t}r(s_{t})}]$, where we use $\mathbb{E}_{\pi}$ to denote the expectation under the distribution induced by $a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim p(\cdot|s_{t},a_{t})$ and $\Delta({\mathcal{A}})$ denotes a probability simplex supported over ${\mathcal{A}}$.

### v2

# RLZero: Direct Policy Inference from Language Without In-Domain Supervision

@@ Abstract
The reward hypothesis states that all goals and purposes can be understood as the maximization of a received scalar reward signal. However, in practice, defining such a reward signal is notoriously difficult, as humans are often unable to predict the optimal behavior corresponding to a reward function. Natural language offers an intuitive alternative for instructing reinforcement learning (RL) agents, yet previous language-conditioned approaches either require costly supervision or test-time training given a language instruction. In this work, we present a new approach that uses a pretrained RL agent trained using only unlabeled, offline interactions—without task-specific supervision or labeled trajectories—to get zero-shot test-time policy inference from arbitrary natural language instructions. We introduce a framework comprising three steps: imagine, project, and imitate. First, the agent imagines a sequence of observations corresponding to the provided language description using video generative models. Next, these imagined observations are projected into the target environment domain. Finally, an agent pretrained in the target environment with unsupervised RL instantly imitates the projected observation sequence through a closed-form solution. To the best of our knowledge, our method, RLZero, is the first approach to show direct language-to-behavior generation abilities on a variety of tasks and environments without any in-domain supervision. We further show that components of RLZero can be used to generate policies zero-shot from cross-embodied videos, such as those available on YouTube, even for complex embodiments like humanoids.
@@ 1 Introduction
Underlying the many successes of RL lies the engineering challenge of reward design. It often takes a skilled expert to engineer a reward function which still might not be fully aligned to the true task specification [9]. Not only does this challenge restrict the scaling of RL agents, but it also makes those agents inaccessible to any user inexperienced in reward design. Even for experts, confidently specifying reward functions is generally infeasible because learning agents have often been found to hack them [38, 4, 18, 31]; i.e. the learned policies for the reward function produce behaviors that do not align with what the human intended. Language is an expressive communication channel for human intent and allows bypassing traditional reward design, but learning a mapping from language to behaviors has historically required collecting and annotating behaviors that correspond to language [26, 33, 54] in the domain where the agent is to be deployed. Scaling this strategy is impractical as it requires labeling the agent’s large space of behaviors with corresponding language descriptions, and repeating the process in every new domain.
@@
Figure 1: RLZero framework of imagine, project, and imitate: A video trajectory is imagined using the text prompt and each frame is projected to agent’s observation space. A closed form solution to imitation learning for BFMs trained with unsupervised RL is used to obtain a policy that mimics the projected video behavior.
This paper is motivated by the vision that a true generalist agent should be able to solve new tasks that are presented to it in a way that is natural for people to specify. Language offers an intuitive channel for task specification, but prior language-based policy learning methods have required significant agent training for specific language commands. In contrast, the objective of this work is to enable generalist agents to translate language commands into behaviors, with no post-command training. The solution we identify can be divided into two parts: (1) being able to communicate the task specification through language commands, and, (2) quickly producing policies for the inferred task. Large-scale multimodal foundation models [79] provide us with the first part of the solution. Trained on large amounts of internet data, they can generate video segments that communicate what performing a task entails. Unfortunately, these video segments can be out-of-distribution of the agent’s domain with different dynamics and embodiments. Thus, this work addresses this challenge by projecting individual frames of generated video to agent observations using semantic similarity [61, 89], thus grounding them in the agent’s observation space.
This presents us with the next part of the problem: How do we quickly infer a policy that solves the communicated task i.e. mimics the grounded imagined trajectories? We utilize the recent success of successor-features-based unsupervised RL methods, sometimes called Behavior Foundation Models (BFMs) [76, 55, 2, 67] for direct policy inference. These models are pretrained in the agent’s domain using task-agnostic prior interactions of arbitrary quality from the environment to encode a wide variety of behaviors. During inference, near optimal policies corresponding to any reward function can be obtained in a zero-shot manner, i.e. in closed form without any gradient updates.
@@
Putting these two parts together, we frame the problem of language-to-policy as a translation from one space, of language, to another space, of behaviors, as specified by successor features. To address this problem, we leverage the capabilities both of video language models to act as a pretrained medium for translating language to states, and of unsupervised RL methods to provide a zero-shot solution for distribution matching. The resulting approach is inspired by the imagination capabilities of humans to picture in their mind possibilities in the real world [63, 64, 59] and then ro rely on past experiences, memories, and abilities to inform their actions. Our framework RLZero (illustrated in figure 1) comprises three main steps: a) Imagine: Imagine trajectories given a language command; b) Project: The frames of imagined trajectories are projected to real observations of the agent; c) Imitate: Leverage a pretrained BFM to directly output a policy that matches the state visitation distribution of the imagined trajectories. Our experiments indicate that RLZero is a promising approach to designing an interpretable link connecting humans to RL agents. We find that RLZero is an effective method on a variety of tasks where reward function design would require an expert reward engineer. We further find that RLZero exhibits successful direct cross-embodiment transfer, the first approach to be able to do so, to the best of our knowledge. This paper’s main contribution is the framework of imagine, project, and zero-shot imitate, which diverges from the prior approach of using VLMs as reward functions—which can be hacked—and instead focuses on zero-shot imitation with unsupervised RL, which admits a unique solution that matches the imagined behavior.
@@ 2 Related Work
Language and Vision Based Control: There has been a rich history of using vision [13, 32, 68, 45] and language [73, 69, 27, 48] as task primitives in RL. While several of these infer a reward function from language instructions and/or VLMs [62, 6, 68, 85, 88, 43] and require training an RL agent using this reward, others have used trajectory-instruction pairs to train the agents with the hope of generalizing to related instructions. A number of recent works have utilized large pretrained foundation models to generate video plans [16, 17, 44] or plans containing textual actions [29, 3, 80]. Some similar methods generate these video plans using diffusion models or energy models [43]. But these methods often require in-domain labeled trajectories to finetune the large models or train these diffusion models [43, 44]. These often rely on APIs [39], inverse dynamics models [16, 44], or goal-conditioned policies trained using text-expert trajectory pairs [17] to convert these plans to actions in the agent’s space. This means that these methods obtain an open-loop sequence of actions that is a translation from the video or textual plan to the agent’s action space. RLZero differs from all of these as: (1) It does not have any test time training or planning and directly obtains an executable closed loop policy, (2) it does not require any in-domain task-dependent supervision like expert labeled trajectories to either train or finetune the video trajectory generation or for policy training. Recent work [49] proposed an approach to generate the video plans without requiring in-domain supervision. But this work still requires training the RL agent for each given task prompt which can be tedious and time consuming.
Zero-shot RL: Zero-shot RL [77] promises the ability to quickly produce optimal policies for any given task defined by a reward function. A wide variety of methods have been developed to achieve zero-shot RL, which are in some ways generalizations of multi-task RL [12]. Most of these works assume a class of tasks where they can produce policies zero-shot. These tasks can be goal-conditioned [34, 19, 1, 65, 47], a linear span of certain state-features [15, 5] or some combination of some skills [20, 21, 56]. Recent works [84, 76, 77, 55, 2] employ a successor measure-based representation learning objective to be able to provide near-optimal policies for arbitrary reward function subject to model capacity constraints. Our work leverages these methods and finds the best reward supported by the representations that will produce the language-conditioned imagined trajectory.
@@ 3 Preliminaries
RLZero uses generative models to imagine trajectories from language prompts and directly produces a policy that results in observations imitating the imagined trajectory. In this section, we introduce the notion of trajectory generation, imitation learning, and unsupervised RL.